# CultureLLM: Incorporating Cultural Differences into Large Language Models

### v1

@@ 6 Conclusion and Limitation
Cultural difference is essential to the prosperity of the world. In this paper, we proposed CultureLLM, a cost-effective solution to fine-tune culture-aware LLMs. We sampled a small number (50) of samples from World Value Survey and then generated augmented data through our novel semantic data augmentation. On $59$ datasets on $9$ cultures, CultureLLM outperformed GPT-3.5 and Gemini with comparable or even better results than GPT-4.
This work has the following limitations. First, due to resource and time constraints, we did not implement CultureLLM on large-scale open-source models. Second, we only adopted classification tasks for evaluation since multilingual generative tasks are expensive for automatic evaluation. Third, for Arabic and Spanish, which are the main spoken languages for many countries, we only select some representative countries to fine-tune for reality factors, and this might also bring bias. Finally, the sample diversity is only in sentence and word levels. In the future, we plan to add more diversities to enrich the generated data.
Acknowledgement
We thank Prof. Diyi Yang from Stanford University for her constructive comments.
Impact Statement
@@ Acknowledgement
Appendix
@@
Appendix A Seed Data from World Values Survey
LABEL:tb-wvs shows the 50 seed samples we used from WVS.
Appendix B Dataset
@@
Table 3: A brief introduction of the $8$ evaluation tasks and $59$ datasets. We list both the name and the size of test sets. For instance, “OffensEval2020(2000) [2020]” denotes that there are 2000 test samples in the dataset OffensEval2020.
Culture
Country & Territory
Task & Dataset
#Sample
Arabic (CultureLLM-Ar)
Middle East
@@
Offensive language detection: OffensEval2020(2000) [2020], OSACT4(1000) [2020], Multi-Platform(1000) [2020], and OSACT5(2541) [2022]. Hate detection: OSACT4(1000) [2020], Multi-Platform(675) [2020], OSACT5(2541) [2022], and OSACT5_finegrained(2541) [2022]. Spam detection: ASHT(1000) [2024]. Vulgar detection: Multi-Platform(675) [2020]
14,973
Bangli (CultureLLM-Bn)
Bangladesh
@@
Offensive language detection: TRAC2020 Task1(1000) [2020], TRAC2020 Task2(1000) [2020], BAD(1000) [2022]. Hate detection: Hate Speech(1000) [2021]. Threat detection: BACD(1000) [2018]. Bias detection: BACD(1000) [2018].
6,000
Chinese (CultureLLM-Zh)
China
@@
Spam detection: CCS(1000) [2019]. Bias detection: CDial-Bias(1000) [2022]. Stance detection: CValues(1712) [2023].
3,712
English (CultureLLM-En)
United States
@@
Offensive language detection: SOLID(1000) [2020]. Hate detection: MLMA(1000) [2019] and HOF(1000) [2017]. Threat detection: CValuesJMT(1000) [2019]. Toxicity detection: MLMA(1000) [2019] and JMT(1000) [2019].
6,000
German (CultureLLM-De)
Germany and parts of Europe
@@
Offensive language detection: GermEval2018(3531) [2018]. Hate detection: IWG_1(469) [2016], IWG_2(469) [2016], HASOC2020(850) [2020], and multilingual-hatecheck(1000) [2022].
6,319
Korean (CultureLLM-Ko)
South Korea
@@
Hate detection: K-MHaS(1000) [2022], hateSpeech(1000) [2020], and HateSpeech2(1000) [2020]. Abusive detection: AbuseEval(1000) [2020], CADD(1000) [2021], and Waseem(1000) [2016].
5,000
Portuguese (CultureLLM-Pt)
Brazil and parts of Latin America
@@
Offensive language detection: OffComBR(1250) [2017], and HateBR(1000) [2022]. Bias detection: ToLD-Br-homophobia(1000) [2020], and ToLD-Br-misogyny(1000) [2020]. Abusive detection: ToLD-Br-insult(1000) [2020].
16,250
Spanish (CultureLLM-Es)
Argentina, Mexico, and parts of Latin America
@@
Offensive language detection: AMI(1000) [2018], MEX-A3T(1000) [2018], and OffendES(1000) [2021]. Hate detection: HatEval 2019(1000) [2019], and HaterNet(1000) [2019]. Bias detection: DETOXIS_stereotype(1000) [2021], and DETOXIS_improper(1000) [2021]. Abusive detection: DETOXIS_abusive(1000) [2021], DETOXIS_mockery(1000) [2021]. Aggressiveness detection: DETOXIS_aggressiveness(1000) [2021]. Stance detection: DETOXIS_stance(1000) [2021].
11,000
Turkish (CultureLLM-Tr)
Turkey
@@
Offensive language detection: SemEval-2020(3528) [2020], offenseCorpus(1000) [2020], offenseKaggle(1000) [2021], and offenseKaggle_2(1000) [2022]. Abusive detection: ATC(1000) [2021]. Spam detection: Turkish Spam(825) [2019]. Fine-grained offensive detection: offenseCorpus(1000) [2020].
10,353
All (CultureLLM-One)
All
All
68,607
B.1 Arabic
@@
OSACT5 [Mubarak et al., 2022] dataset consists of 12,698 Arabic tweets collected between June 2016 and November 2017, labeled for offensiveness and fine-grained hate speech types using emojis commonly found in offensive communications, providing a resource for offensive and hate speech detection and classification tasks. The dataset consists of three subtasks: offensiveness detection, hate speech detection, and fine-grained hate speech detection. We utilized 2,541 data samples for each of these tasks.
ASHT [Kaddoura and Henno, 2024] dataset contains 132,421 Arabic tweets collected from Twitter, classified as either ham (non-spam) or spam, providing a valuable resource for researchers in Arabic natural language processing (NLP) and serving as a benchmark for research in Arabic NLP, cybersecurity, data science, and social network analysis. We utilized a subset of 1,000 data samples for the spam detection section.
B.2 Bengali
@@ B.9 Turkish
Appendix C Prompt Setting
@@
Appendix D Details Experimental Results

### v2

@@ 6 Conclusion and Limitation
Cultural difference is essential to the prosperity of the world. In this paper, we proposed CultureLLM, a cost-effective solution to fine-tune culture-aware LLMs. We sampled a small number (50) of samples from the World Value Survey and then generated augmented data through our novel semantic data augmentation. On $59$ datasets on $9$ cultures, CultureLLM outperformed GPT-3.5 and Gemini with comparable or even better results than GPT-4.
This work has the following limitations. First, due to resource and time constraints, we did not implement CultureLLM on large-scale open-source models. Second, we only adopted classification tasks for evaluation since multilingual generative tasks are expensive for automatic evaluation. Finally, the sample diversity is only in sentence and word levels. In the future, we plan to add more diversities to enrich the generated data.
Disclaimer
@@
Appendix A Discussion on the Relationship between Culture and Language
We strongly agree that language is not equal to, but only a part of culture. But using language to study culture is possible due to the following aspects:
  1. 1.
    Existing literature on culture understanding shows that culture boundaries are fluid, dynamic and uncertain. Delanoy emphasizes that cultures are not homogeneous or static entities but are fluid and dynamic. He critiques essentialist views that rigidly define cultural boundaries and instead promotes a more nuanced understanding that considers the intersections of various cultural factors, such as ethnicity, language, religion, and socio-economic conditions [Delanoy, 2020a]. Appadurai also discusses the fluidity of cultural boundaries and the creation of new cultural forms [Appadurai, 1996]. Cultural boundaries can be geographical regions, language, religion and so on. Based on above statements, using language as cultural boundaries is reasonable.
  2. 2.
    Existing NLP works on culture also leverage labguage as culture boundaries. Naous et al. [2023] focuses on Arabic and English culture.  Wang et al. [2023d] focuses on 8 different cultures: English, Chinese, French, Russian, German, Arabic, Japanese and Korean. Liu et al. [2023b] also use language to split different cultures. The authors work on English, German, Russian, Bengali, Chinese, and Indonesian culture. Myung et al. [2024] is a hand-crafted benchmark for evaluate diverse cultures. They also use languages as culture boundaries.
  3. 3.
    Most downstream benchmarks are classified via language and we cannot get more fine-grained perspectives. For example, if we want to evaluate the performance of Arabic model, we can find benchmarks in Arabic culture. But if we use regions as cultural boundaries, we can’t find benchmarks in Morocco and Jordan cultures.
Appendix B Details on Data
B.1 World Values Survey and seed data
The following is the basic information of WVS and Table 3 shows the 50 seed data.
World Values Survey (WVS) is a global research initiative dedicated to rigorously examining the diverse array of social, political, economic, religious, and cultural values held by people worldwide. Its overarching aim is to analyze the impact that shifts in these values, whether stable or evolving, exert on the development of countries and societies across various dimensions. Originating from the European Values Study, the project was initiated in 1981 by Professor Ronald Inglehart and his team from the University of Michigan (USA), who served as its first President until 2013. Operating in over 120 societies worldwide, the WVS conducts a comprehensive comparative social survey every five years, serving as its primary research tool. Its expansive coverage, encompassing diverse geographies and topics, coupled with the open access to survey data and findings, has solidified the WVS as one of the most reputable and extensively utilized cross-national surveys in the realm of social sciences. Currently, it is the most extensive non-commercial investigation into human beliefs and values over time, making significant contributions to empirical understanding on a global scale.
The preferred sampling method for the World Values Survey involves selecting a full probability sample from the population aged 18 years and older. Typically, this requires access to a comprehensive list or registry of all households or voters within the country. However, acknowledging the potential financial constraints associated with full probability samples, the WVSA permits the utilization of a nationally representative random sample through multi-stage territorial stratified selection. Each national team tailors its sampling model to accommodate the unique characteristics of their country, including geographical and administrative divisions, urban and rural population sizes, available statistical data, and adherence to WVSA methodological standards. For a country to employ the WVS-7 sample model, it must meet specific criteria: - WVS surveys are required to cover all residents (not just citizens) in a country in the age of 18 years older and older; - PI’s can lower the minimum age limit as long as the minimum required sample size for the 18+ population is achieved; - Obtained sample should be representative, i.e. should reflect the main distributions observed in the country population (gender; age groups; urban/rural population etc.).
We use WVS-7 (2017-2022) dataset V5.0 which started in mid-2017 and following a 1-year postponement due to the Covid-pandemic, was finally closed on December 31, 2021. It covers 80 countries. Samples must be representative of all people in the age 18 and older residing within private households in each country, regardless of their nationality, citizenship or language. The minimum sample size - i.e. the number of completed interviews which are included into the national data-set in most of countries is 1200. Countries with greater population size and diversity apply samples of N=1500 to N=5000. Countries with the population below 2 million people apply samples of N=1000.
Regards on the select criteria for seed data, we select both low-resource cultures and high-resource cultures for comparison. WVS has 294 questions in total, and we manually rewrite 50 questions out of them into QA format. The selection criterion is to ask several cultural experts and LLMs and we selected where both of them have the same confidence.
The rewrite process is to manually rewrite the questions (multi-choices format) from WVS into QA format. For example, the original question "Do your agree with One of my main goals in life has been to make my parents proud?" can be rewritten into "Give me the answer from 1 to 4: Do you agree with One of my main goals in life has been to make my parents proud? 1. Strongly agree 2. agree 3. Disagree 4. Strongly disagree. You can only choose one option." We select seed data that contains each topic mentioned in WVS.
Before consolidating the average responses from representative countries, we undertake a meticulous manual review of their answers to ensure consistency on identical questions. In instances where the responses from various countries significantly diverge, it highlights the necessity for us to consider fine-tuning distinct models in future endeavors. This careful preliminary check ensures that our analysis remains robust and that any future model adaptations are informed by a clear understanding of cultural variances.
@@ B.1 World Values Survey and seed data
Table 3: Seed data from World Values Survey. The same questions can be paired with opinions from different cultures.
Topic Question
SOCIAL VALUES Do you agree with One of my main goals in life has been to make my parents proud?
Do you agree with When a mother works for pay, the children suffer?
Do you agree with On the whole, men make better political leaders than women do?
Do you agree with A university education is more important for a boy than for a girl?
Do you agree with On the whole, men make better business executives than women do?
Do you agree with Being a housewife is just as fulfilling as working for pay?
Do you agree with When jobs are scarce, men should have more right to a job than women?
Do you agree with When jobs are scarce, employers should give priority to people of this country over immigrants?
Do you agree with If a woman earns more money than her husband, it’s almost certain to cause problems?
Do you agree with Homosexual couples are as good parents as other couples?
Do you agree with It is a duty towards society to have children?
Do you agree with Adult children have the duty to provide long-term care for their parents?
Do you agree with People who don’t work turn lazy?
Do you agree with Work is a duty towards society?
Do you agree with Work should always come first, even if it means less spare time?
MIGRATION In terms of the effects of immigration on the development of your country, do you agree with Fills important jobs vacancies?
In terms of the effects of immigration on the development of your country, do you agree with Strengthens cultural diversity?
In terms of the effects of immigration on the development of your country, do you agree with Increases the crime rate?
In terms of the effects of immigration on the development of your country, do you agree with Gives asylum to political refugees who are persecuted elsewhere?
In terms of the effects of immigration on the development of your country, do you agree with Increases the risks of terrorism?
In terms of the effects of immigration on the development of your country, do you agree with Offers people from poor countries a better living?
In terms of the effects of immigration on the development of your country, do you agree with Increases unemployment?
In terms of the effects of immigration on the development of your country, do you agree with Leads to social conflict?
SECURITY How frequently do the following things occur in your neighborhood: Robberies?
How frequently do the following things occur in your neighborhood: Alcohol consumption in the streets?
How frequently do the following things occur in your neighborhood: Police or military interfere with people’s private life?
How frequently do the following things occur in your neighborhood: Racist behavior?
How frequently do the following things occur in your neighborhood: Drug sale in streets?
How frequently do the following things occur in your neighborhood: Street violence and fights?
How frequently do the following things occur in your neighborhood: Sexual harassment?
SCIENCE Do you agree with Science and technology are making our lives healthier, easier, and more comfortable.?
Do you agree with Because of science and technology, there will be more opportunities for the next generation.?
Do you agree with We depend too much on science and not enough on faith.?
Do you agree with One of the bad effects of science is that it breaks down people’s ideas of right and wrong.?
Do you agree with It is not important for me to know about science in my daily life.?
RELI GION Do you agree with Whenever science and religion conflict, religion is always right?
Do you agree with The only acceptable religion is my religion.?
ETHICS Do you think that the your country’s government should or should not have the right to do the following: Keep people under video surveillance in public areas?
Do you think that the your country’s government should or should not have the right to do the following: Monitor all e-mails and any other information exchanged on the Internet?
Do you think that the your country’s government should or should not have the right to do the following: Collect information about anyone living in this country without their knowledge?
POLITICAL In your view, how often do the following things occur in this country’s elections: Votes are counted fairly?
In your view, how often do the following things occur in this country’s elections: Opposition candidates are prevented from running?
In your view, how often do the following things occur in this country’s elections: TV news favors the governing party?
In your view, how often do the following things occur in this country’s elections: Voters are bribed?
In your view, how often do the following things occur in this country’s elections: Journalists provide fair coverage of elections?
In your view, how often do the following things occur in this country’s elections: Election officials are fair?
In your view, how often do the following things occur in this country’s elections: Rich people buy elections?
In your view, how often do the following things occur in this country’s elections: Voters are threatened with violence at the polls?
In your view, how often do the following things occur in this country’s elections: Voters are offered a genuine choice in the elections?
In your view, how often do the following things occur in this country’s elections: Women have equal opportunities to run the office
B.2 Details on experimental datasets
@@
Table 4: A brief introduction of the $8$ evaluation tasks and $59$ datasets. We list both the name and the size of test sets. For instance, “OffensEval2020(2000) [Zampieri et al., 2020]” denotes that there are 2000 test samples in the dataset OffensEval2020.
Culture Country & Territory Task & Dataset #Sample
@@
Arabic (CultureLLM-Ar) Middle East Offensive language detection: OffensEval2020(2000) [Zampieri et al., 2020], OSACT4(1000) [Husain, 2020], Multi-Platform(1000) [Chowdhury et al., 2020], and OSACT5(2541) [Mubarak et al., 2022]. Hate detection: OSACT4(1000) [Husain, 2020], Multi-Platform(675) [Chowdhury et al., 2020], OSACT5(2541) [Mubarak et al., 2022], and OSACT5_finegrained(2541) [Mubarak et al., 2022]. Spam detection: ASHT(1000) [Kaddoura and Henno, 2024]. Vulgar detection: Multi-Platform(675) [Chowdhury et al., 2020] 14,973
@@
Bangli (CultureLLM-Bn) Bangladesh Offensive language detection: TRAC2020 Task1(1000) [Bhattacharya et al., 2020], TRAC2020 Task2(1000) [Bhattacharya et al., 2020], BAD(1000) [Sharif and Hoque, 2022]. Hate detection: Hate Speech(1000) [Romim et al., 2021]. Threat detection: BACD(1000) [aimansnigdha, 2018]. Bias detection: BACD(1000) [aimansnigdha, 2018]. 6,000
@@
Chinese (CultureLLM-Zh) China Spam detection: CCS(1000) [Jiang et al., 2019]. Bias detection: CDial-Bias(1000) [Zhou et al., 2022]. Stance detection: CValues(1712) [Xu et al., 2023]. 3,712
@@
English (CultureLLM-En) United States Offensive language detection: SOLID(1000) [Rosenthal et al., 2020]. Hate detection: MLMA(1000) [Ousidhoum et al., 2019] and HOF(1000) [Davidson et al., 2017]. Threat detection: CValuesJMT(1000) [Kaggle, 2019]. Toxicity detection: MLMA(1000) [Ousidhoum et al., 2019] and JMT(1000) [Kaggle, 2019]. 6,000
@@
German (CultureLLM-De) Germany and parts of Europe Offensive language detection: GermEval2018(3531) [Wiegand et al., 2018]. Hate detection: IWG_1(469) [Ross et al., 2016], IWG_2(469) [Ross et al., 2016], HASOC2020(850) [HASOC, 2020], and multilingual-hatecheck(1000) [Röttger et al., 2022]. 6,319
@@
Korean (CultureLLM-Ko) South Korea Hate detection: K-MHaS(1000) [Lee et al., 2022], hateSpeech(1000) [Moon et al., 2020], and HateSpeech2(1000) [daanVeer, 2020]. Abusive detection: AbuseEval(1000) [Caselli et al., 2020], CADD(1000) [Song et al., 2021], and Waseem(1000) [Waseem and Hovy, 2016]. 5,000
@@
Portuguese (CultureLLM-Pt) Brazil and parts of Latin America Offensive language detection: OffComBR(1250) [de Pelle and Moreira, 2017], and HateBR(1000) [Vargas et al., 2022]. Bias detection: ToLD-Br-homophobia(1000) [Leite et al., 2020], and ToLD-Br-misogyny(1000) [Leite et al., 2020]. Abusive detection: ToLD-Br-insult(1000) [Leite et al., 2020]. 16,250
@@
Spanish (CultureLLM-Es) Argentina, Mexico, and parts of Latin America Offensive language detection: AMI(1000) [Fersini et al., 2018], MEX-A3T(1000) [Álvarez-Carmona et al., 2018], and OffendES(1000) [Plaza-del Arco et al., 2021]. Hate detection: HatEval 2019(1000) [Basile et al., 2019], and HaterNet(1000) [Pereira-Kohatsu et al., 2019]. Bias detection: DETOXIS_stereotype(1000) [de Paula and Schlicht, 2021], and DETOXIS_improper(1000) [de Paula and Schlicht, 2021]. Abusive detection: DETOXIS_abusive(1000) [de Paula and Schlicht, 2021], DETOXIS_mockery(1000) [de Paula and Schlicht, 2021]. Aggressiveness detection: DETOXIS_aggressiveness(1000) [de Paula and Schlicht, 2021]. Stance detection: DETOXIS_stance(1000) [de Paula and Schlicht, 2021]. 11,000
@@
Turkish (CultureLLM-Tr) Turkey Offensive language detection: SemEval-2020(3528) [Zampieri et al., 2020], offenseCorpus(1000) [Çöltekin, 2020], offenseKaggle(1000) [Kaggle, 2021], and offenseKaggle_2(1000) [Kaggle, 2022]. Abusive detection: ATC(1000) [Karayiğit et al., 2021]. Spam detection: Turkish Spam(825) [mis, 2019]. Fine-grained offensive detection: offenseCorpus(1000) [Çöltekin, 2020]. 10,353
All (CultureLLM-One) All All 68,607
B.2.1 Arabic
@@
OSACT5 [Mubarak et al., 2022] dataset consists of 12,698 Arabic tweets collected between June 2016 and November 2017, labeled for offensiveness and fine-grained hate speech types using emojis commonly found in offensive communications, providing a resource for offensive and hate speech detection and classification tasks. The dataset consists of three subtasks: offensiveness detection, hate speech detection, and fine-grained hate speech detection. We used 2,541 samples for each of these tasks.
ASHT [Kaddoura and Henno, 2024] dataset contains 132,421 Arabic tweets collected from Twitter, classified as either ham (non-spam) or spam, providing a valuable resource for researchers in Arabic natural language processing (NLP) and serving as a benchmark for research in Arabic NLP, cybersecurity, data science, and social network analysis. We used a subset of 1,000 samples for the spam detection section.
B.2.2 Bengali
@@ B.2.9 Turkish
Appendix C Evaluation Metrics and Prompts
C.1 Evaluation metrics
C.1.1 Perplexity
C.1.2 Diversity Gain
where $f$ is the feature extractor, and we use OpenAI Embedding API text-embedding-ada-002 to extract features.
C.2 Prompts for evaluation
Table 5: Prompts for evaluation.
Task Prompt
offensive_detect if the following sentence is offensive, just say \"OFF\", otherwise, say \"NOT_OFF\" without explanation:
hate_detect if the following sentence has hate speech, just say \"HS\", otherwise, say \"NOT_HS\" without explanation:
vulgar_detect if the following sentence has vulgar speech, just say \"Vulgar\", otherwise, say \"NOT_Vulgar\" without explanation:
spam_detect if the following sentence is spam tweet, just say \"Spam\", otherwise, say \"NOT_Spam\" without explanation:
stereotype_detect if the following sentence has stereotype speech, just say \"1\", otherwise, say \"0\" without explanation:
mockery_detect if the following sentence has mockery speech, just say \"1\", otherwise, say \"0\" without explanation:
insult_detect if the following sentence has insult speech, just say \"1\", otherwise, say \"0\" without explanation:
improper_detect if the following sentence has improper speech, just say \"1\", otherwise, say \"0\" without explanation:
aggressiveness_detect if the following sentence has aggressiveness speech, just say \"1\", otherwise, say \"0\" without explanation:
toxicity_detect if the following sentence has toxicity speech, just say \"1\", otherwise, say \"0\" without explanation:
negative_stance_detect if the following sentence has negative stance speech, just say \"1\", otherwise, say \"0\" without explanation:
homophobia_detect if the following sentence has homophobia speech, just say \"1\", otherwise, say \"0\" without explanation:
racism_detect if the following sentence has racism speech, just say \"1\", otherwise, say \"0\" without explanation:
misogyny_detect if the following sentence has misogyny speech, just say \"1\", otherwise, say \"0\" without explanation:
threat_detect if the following sentence has threat speech, just say \"1\", otherwise, say \"0\" without explanation:
bias_on_gender_detect if the following speech expressing bias on gender, just say \"1\", otherwise, say \"0\" without explanation:
hostility_directness_detect if the following speech expressing hostility directness, just say \"1\", otherwise, say \"0\" without explanation:
hate_offens_detect if the following sentence contains hate speech, just say \"0\", else if contains offensive language, say \"1\", otherwise, say \"2\" without explanation:
hate_detect_fine-grained if the following sentence doesn’t have hate speech, just say \"NOT_HS\", otherwise, label the hate speech with \"HS1\"(Race), \"HS2\"(Religion), \"HS3\"(Ideology), \"HS4\"(Disability), \"HS5\"(Social Class), \"HS6\"(Gender) without explanation:
offensive_detect_finegrained if the following sentence doesn’t have offensive speech, just say \"non\", otherwise, label the offensive speech with \"prof\"(profanity, or non-targeted offense), \"grp\"(offense towards a group), \"indv\"(offense towards an individual), \"oth\"(ffense towards an other (non-human) entity, often an event or organization) without explanation:
For evaluation, we input a system prompt to induce LLM to specific culture. For example, we add “You are an Arabic chatbot that knows Arabic very well.” as system prompt for evaluation on Arabic understanding. As for the prompt for evaluation on different tasks, the detailed prompts are shown in Table 5.
Prompts for open-ended tasks:
@@ C.2 Prompts for evaluation
C.3 Prompts for RAG
Table 6: Information retrieved via RAG
Culture Information
Arabic Arab culture is the culture of the Arabs, from the Atlantic Ocean in the west to the Arabian Sea in the east, in a region of the Middle East and North Africa known as the Arab world. The various religions the Arabs have adopted throughout their history and the various empires and kingdoms that have ruled and took lead of the civilization have contributed to the ethnogenesis and formation of modern Arab culture. Language, literature, gastronomy, art, architecture, music, spirituality, philosophy and mysticism are all part of the cultural heritage of the Arabs.
Bengali The culture of Bengal defines the cultural heritage of the Bengali people native to eastern regions of the Indian subcontinent, mainly what is today Bangladesh and the Indian states of West Bengal and Tripura, where they form the dominant ethnolinguistic group and the Bengali language is the official and primary language. Bengal has a recorded history of 1,400 years. The Bengalis are dominant ethnolinguistic group. The Bengal region has been a historical melting point, blending indigenous traditions with cosmopolitan influences from pan-Indian subcontinental empires. Dhaka (Dacca) became the capital of Mughal Bengal (Bengal Subah) and the commercial (financial) capital (1610-1757) of Mughal India. Dhaka is the largest and richest Bengali (Bangali) mega city in the world and also the 3rd largest and richest mega city in (Indian sub continent) after Mumbai (Bombay or MMR) and Delhi (NCR). Dhaka is a Beta Global City (Moderate Economic Centre). As a part of the Bengal Presidency, Bengal also hosted the region’s most advanced political and cultural centers during British rule.
Chinese Chinese culture is one of the world’s oldest cultures, originating thousands of years ago. The culture prevails across a large geographical region in East Asia with Sinosphere in whole and is extremely diverse, with customs and traditions varying greatly between counties, provinces, cities, towns. The terms ’China’ and the geographical landmass of ’China’ have shifted across the centuries, before the name ’China’ became commonplace in modernity. Chinese civilization is historically considered a dominant culture of East Asia. With China being one of the earliest ancient civilizations, Chinese culture exerts profound influence on the philosophy, virtue, etiquette, and traditions of Asia. Chinese characters, ceramics, architecture, music, dance, literature, martial arts, cuisine, arts, philosophy, etiquette, religion, politics, and history have had global influence, while its traditions and festivals are celebrated, instilled, and practiced by people around the world.
English The culture of the United States of America, also referred to as American culture, encompasses various social behaviors, institutions, and norms in the United States, including forms of speech, literature, music, visual arts, performing arts, food, sports, religion, law, technology as well as other customs, beliefs, and forms of knowledge. American culture has been shaped by the history of the United States, its geography, and various internal and external forces and migrations. Several historical ethnicities make up American culture: Yankee (Anglo-American), Tejano, Louisiana French (Cajun, Creole), Pennsylvania Dutch (Fancy Dutch, Amish), New York Dutch, Texas German, Alaskan Russian, Puerto Rican, Hawaiian.
Germany The culture of Germany has been shaped by major intellectual and popular currents in Europe, both religious and secular. German culture originated with the Germanic tribes, the earliest evidence of Germanic culture dates to the Jastorf culture in Northern Germany and Denmark. Contact with Germanic tribes were described by various Greco-Roman authors. The first extensive writing done on Germanic culture can be seen during the Roman Imperial Period with Germania by Tacitus.
Korean The contemporary culture of South Korea developed from the traditional culture of Korea which was prevalent in the early Korean nomadic tribes. By maintaining thousands of years of ancient Korean culture, with influence from ancient Chinese culture, South Korea split on its own path of cultural development away from North Korean culture since the division of Korea in 1948. The industrialization, urbanization and westernization of South Korea, especially Seoul, have brought many changes to the way Korean people live. Changing economics and lifestyles have led to urbanization—a concentration of population in major cities (and depopulation of the rural countryside), with multi-generational households separating into nuclear family living arrangements. Today, many cultural elements from South Korea, especially popular culture, have spread across the globe and have become some of the most prominent cultural forces in the world.
Portuguese The culture of Portugal is a very rich result of a complex flow of many different civilizations during the past millennia. From prehistoric cultures, to its Pre-Roman civilizations (such as the Lusitanians, the Gallaeci, the Celtici, and the Cynetes, amongst others), passing through its contacts with the Phoenician-Carthaginian world, the Roman period (see Hispania, Lusitania and Gallaecia), the Germanic invasions of the Suebi, Buri (see Kingdom of the Suebi) and Visigoths (see Visigothic Kingdom), Viking incursions, Sephardic Jewish settlement, and finally, the Moorish Umayyad invasion of Hispania and the subsequent expulsion, during the Reconquista, all have made an imprint on the country’s culture and history. The name of Portugal itself reveals much of the country’s early history, stemming from the Roman name Portus Cale, a Latin name meaning \"Port of Cale\" (Cale likely is a word of Celtic origin - Cailleach-Bheur her other name; the Mother goddess of the Celtic people as in Calais, Caledonia, Beira. She was the one who, with a hammer created mountains and valleys; the one who hid in stones and trees - Mother nature), later transformed into Portucale, and finally into Portugal, which emerged as a county of the Kingdom of León see County of Portugal) and became an independent kingdom in 1139. During the 15th and 16th centuries, Portugal was a major economic, political, and cultural power, its global empire stretching from the Americas, to Africa, and various regions of Asia and Oceania. Portugal, as a country with a long history, is home to several ancient architectural structures, as well as typical art, furniture and literary collections mirroring and chronicling the events that shaped the country and its peoples. It has a large number of cultural landmarks ranging from museums to ancient church buildings to medieval castles, which testify its rich national cultural heritage. Portugal is home to fifteen UNESCO World Heritage Sites, ranking it 8th in Europe and 17th in the world.
Spanish The culture of Spain is influenced by its Western origin, its interaction with other cultures in Europe, its historically Catholic religious tradition, and the varied national and regional identities within the country. It encompasses literature, music, visual arts, cuisine as well as contemporary customs, beliefs, institutions, and social norms. Beyond Spain, Spanish culture is the foundation of most of Latin American cultures and the Filipino culture. The ancient peoples of Spain included Tartessians, Celts, Iberians, Celtiberians, Phoenicians as well as Greek colonies. Spain largely came under the rule of Carthage and was then entirely conquered by Rome, becoming a province of the Roman empire. The name of Spain derives from the Latin term Hispania, itself a name of Punic origin. In the areas of language and religion, the ancient Romans left a lasting cultural, legal and administrative legacy in the Spanish history. The subsequent course of Spanish history added new elements to the country’s culture and traditions. The Visgoths established a united Hispania and kept the Latin and Christian legacy in Spain between the fall of the Roman Empire and the Early Middle Ages. Muslim influences played a significant role during the Middle Ages in the areas conquered by the Umayyads. However, these influences were not completely assimilated into the Spanish culture, leading to conflicts and ultimately to the Christian Reconquista that would largely shape the culture of the country.
Turkish The culture of Turkey (Turkish: Türkiye kültürü) or the Turkish culture (Türk kültürü) combines a heavily diverse and heterogeneous set of elements that have been derived from the various cultures of the Eastern European, Eastern Mediterranean, Caucasian, Middle Eastern and Central Asian traditions. Many of these traditions were initially brought together by the Ottoman Empire, a multi-ethnic and multi-religious state spanning across Southern Europe, Eastern Europe, the Middle East and North Africa. During the early years of the Republic of Turkey, established after the collapse of the Ottoman Empire, the government invested large sums of resources into fine arts such as architecture and sculpture, and other artistic fields around the country in-line with the newly implemented reformist and West-leaning policies. This was done as part of a process of modernization, westernization, and of creating and outlining a new Turkish cultural identity, rather than the previously established and depicted Ottoman identity.
The detailed information retrieved via RAG is shown in Table 6.
Appendix D Experimental Results